OpenStudy (anonymous):

solve for x x-8/3=12

OpenStudy (mertsj):

Add 8/3 to both sides.

OpenStudy (anonymous):

Answer: 28?

OpenStudy (mertsj):

no

OpenStudy (mertsj):

12+8/3 is NOT 28

OpenStudy (anonymous):

Lol sorry :T hold on

OpenStudy (anonymous):

i got 14.6

OpenStudy (mertsj):

Well, if you want the exact answer, it is 14 2/3

OpenStudy (mertsj):

Or you could put 44/3

OpenStudy (anonymous):

I think it would of been better if you helped me out instead of giving me the answer but thanks

OpenStudy (mertsj):

No good deed goes unpunished.

pooja195 (pooja195):

All you do is add 8/3 to both sides, so 12 + 8/3, or 2 2/3, so x = 14 2/3.

OpenStudy (anonymous):

maybe it is \(\frac{x-3}{8}=12\)

OpenStudy (anonymous):

in which case you multiply both sides by \(8\) to get \(x-3=96\) and then add \(3\) to both sides